OpenStudy (anonymous):

Gary has been given the task of purchasing chips and pretzels for a big backyard picnic. He's been given $30 to spend on the snacks. When he gets to the store, he finds that each bag of chips, x, costs about $2 and each bag of pretzels, y, costs about $3. b) Use whatever method you wish to find 3 possible combinations of bags of chips and pretzels that will cost exactly $30. Even if graphing is not your solution method, explain where these solution points would be found on the graph of the inequality.

OpenStudy (nathan917):

Well lets just say 3/30 = 10 and 2/30=15. so x=15 and y=10

OpenStudy (anonymous):

is that the answer? x=15 and y =10

OpenStudy (nathan917):

2x + 3y = 30 2x = 30 – 3y x = (30 – 3y)/2 = 15 – 3y/2 value of y is multiple of 2 if y = 2 x = 15 – (3 × 2)/2 = 15 – 3 = 12 if y = 4 x = 15 – (3 × 4)/2 = 15 – 6 = 9 if y = 6, x = 6 y = 8, x = 3 y = 10 x = 0 here is the whole question.

OpenStudy (nathan917):

http://malaysia.answers.yahoo.com/question/index?qid=20130402054433AADRusd Which you asked